OpenStudy (amistre64):

rewrite: \[\sum x^2\] : as a function ....

OpenStudy (anonymous):

\[\int{x^2dx}\]

OpenStudy (amistre64):

lol ... not quite what i had in mind but..

OpenStudy (amistre64):

one is discrete, the other you presented is continuous

OpenStudy (amistre64):

at least in my head that is

OpenStudy (anonymous):

\[y = \frac{x^3}{3}\]

OpenStudy (anonymous):

\[+ C\]

OpenStudy (amistre64):

\[\sum x = \frac{x(x+1)}{2}\] \[\sum x^2 = \frac{x(x+1)(2x+1)}{6}\] if i recall correctly

OpenStudy (dumbcow):

yes thats it

OpenStudy (anonymous):

Thought you wanted a riemann sum or something :-P